# Nonlocal Models in Biology and Life Sciences: Sources, Developments, and Applications

### v1

@@ 2.2 Nonlocal interactions in a single species population
Without loss of our generality, we assume that $u$ is a dimensionless population density, and it is defined in the one-dimensional spatial domain $\Omega=[-L,L]\subset\mathbb{R}$, $L>0$. In the considered case $g(u)=u(1-u)$, the population has the reproduction term, which is proportional to its density and available resources $(1-u)$ volpert2009 ; banerjee2022review . For the local interaction, the population’s consumption rate at any given spatial point $x\in\Omega$ is proportional to its population’s density at the same spatial point with the total amount of available resources unity. In general, the individual located at a spatial point $x$ accesses the resources at its location and some in its neighbourhood locations furter1989 ; banerjee2022review ; billingham2020 ; pal2018bmb ; tian2021 ; banerjee2017ec . We assume that the individuals located at the spatial point $y$ access the resources of the point $x$ with the probability density function $\phi(x-y)$. In this case, the RD equation (2) modified into an integro-differential equation
@@
with non-negative initial conditions and periodic boundary conditions over the domain $[-L,L]$. Here $\phi\ast u$ is the convolution of $u$, and we use `$\ast$` ∗’ as the convolution operator for the rest of this manuscript. The convolution $\phi\ast u$ in one spatial dimension is defined by
@@
The nonlocal model (3) has two homogeneous stationary solutions $u=0$ and $u=1$. In the case of the local model, the solution $u=1$ is stable under heterogeneous perturbations, but it may not be stable for the nonlocal model volpert2009 . A brief analysis of this single species nonlocal model is given in the next section.
@@ 2.3 Nonlocal interactions in time and space
Researchers have studied nonlocal RD models for the triangular, parabolic kernel functions to account for different biological phenomena dornelas2021 ; pal2019mbe ; segal2013 ; pal2020amm ; sherratt2016 ; merchant2011 ; pal2019ijbc . These kernel functions are known as the truncated kernel. Such non-truncated kernel functions have been the subject of intensive studies. For instance, Britton introduced an advanced version of the single species model (3) as:
@@
where $\alpha u$ ($\alpha>0$) represents the local aggregation and $-(1+\alpha)\psi\ast u$ is the disadvantage in the global population as the resources may go into depletion of being high population density britton1989 . The author studied the nonlocal model with the Laplacian kernel function sherratt2016 . Sometimes, species take time to move; in this case, the correct average population density is a spatio-temporal weighted average toward the current time and position. Taking this spatio-temporal factor into account, Britton introduced a space-time nonlocality in the single species model (4) as britton1990 :
@@
Likewise, the spatial kernel function $\Phi$ in (3), the spatio-temporal kernel function $\psi$ satisfies some conditions mentioned in britton1990 ; pal2020mbe . In addition, these types of nonlocal models with non-truncated kernels can be converted into coupled RD equations britton1989 ; britton1990 . A brief formulation of this conversion and its applications are given in the upcoming section.
@@ 2.4 Nonlocal interactions in cell biology and genomics
In cell biology, nonlocal interactions are interactions between cellular components that extend beyond immediate neighbours and may involve long-distance signalling or communication. These interactions are critical in cellular processes and behaviours such as cell migration, pattern formation, and collective cell responses. For example, tumors are abnormal masses of tissue consisting of cells with different phenotypes (e.g., cancer stem cells) al2003 ; bonnet1997 ; dick2003 , and they are heterogeneously distributed in the body. Motivated by the paradoxical findings of an agent-based model enderling2009 , Hillen et al. have proposed a mathematical model hillen2013 :
@@
Gene silencing and RNA interference are exciting fields of study that have promising implications in health and industry yang2007 . For instance, a synthetic RNA molecule is designed to selectively target and mute the production of one particular gene in small interfering RNA technology. When it enters cells, it attaches itself to the target mRNA, causing it to be degraded and stopping the creation of new proteins. There are some more key technologies and models associated with gene silencing and RNA interference: short hairpin RNA lambeth2013 , micro RNA macfarlane2010 , CRISPR/Cas9 technology redman2016 , Antisense Oligonucleotides rinaldi2018 , Dicer Knockout models guo2020 , etc. It is essential to model genetic switches and genetic networks in order to comprehend the complex regulatory processes that exist within cells. The dynamics of genetic switches and networks have been simulated and examined using a variety of computational and mathematical models. Some common approaches are Boolean networks, ordinary differential equations, stochastic models, agent-based models, and machine learning approaches. Furthermore, nonlocal models are important in various areas beyond DNA, RNA, and protein modelling. The use of microarray technology for placing the full human genome on chips is one such application.
@@
A multi-scale modelling technique is required to represent a cell as a composite structure incorporating microtubules and organelles. Several models have been created to simulate these components’ interactions and dynamics, such as continuum mechanics models singh2020 ; singh2020iw ; singh2020ic . Understanding the methods by which a cell develops, duplicates its components, and divides into two daughter cells is fundamental to the study of cell cycles. Modelling cell cycles with fractional derivatives presents a mathematical technique that reflects the nonlocal and memory-dependent character of cellular dynamics. In addition, rheology and hysteresis are more frequently linked with materials and physical systems, and they may also be used to describe some aspects of cellular activity, including cell mechanics and cycles melnik2009 . Understanding the interaction of grain shape, material characteristics, and nonlocal rheology is critical for forecasting and regulating the behaviour of dense granular flows fazelpour2022 .
@@
Different types of biological rhythms occur in various physiological and behavioural processes in living organisms, and they are essential for maintaining homeostasis and optimizing physiological functions in organisms. In addition, angiogenesis is the physiological process that results in the formation of new blood vessels from pre-existing vessels carmeliet2005 . This complex mechanism is required for a variety of physiological activities, including embryonic development and wound healing. However, angiogenesis has been linked to a number of pathological disorders, including cancer and inflammatory illnesses. Furthermore, biological cells are complex structures with distinct rheological properties, and understanding their mechanical responses is critical for a variety of applications, including disease diagnostics, medication delivery, and tissue creation. The study of these rhythms provides insights into how living organisms adapt to their environments and regulate various processes over time.
Nonlocal interactions on networks in biology refer to interactions between nodes in a network (such as genes, proteins, or cells) that extend beyond their direct neighbours or immediate connections. These interactions serve as essential for understanding the processes and characteristics of complex biological systems in a network. Furthermore, these interactions on networks are relevant in various biological systems, including signalling networks, ecological food webs, and neuronal networks.
Neurons in the brain communicate and interact through synapses, which can involve long-range connections. Nonlocal models have been employed to describe neurons’ collective activity and synchronization in neural networks. In addition, Alzheimer’s disease (AD) is also growing fast in the current days, and researchers have been trying to understand the mechanisms behind its propagation. Many factors are involved in AD progression, but two protein families, amyloid-beta and tau proteins, are known to be the main contributors matthaus2009 ; bressloff2014 . In thompson2020 , authors have considered the heterodimer model to incorporate the interaction between these two proteins, each consisting of healthy and toxic densities. Also, they derived the network model corresponding to the PDE model for applying it to AD progression in the brain connectome. Authors in pal2022sr modified the heterodimer model by introducing nonlocal interactions into the model. They studied the following nonlocal network model
$\frac{du_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}u_{k}+u_{j}(a_{0}-a_{1}u_{j})-\frac{a_{2}u_{j}}{1+c_{u}u_{j}}\Phi_{j}\ast\widetilde{u_{j}},$ (7a)
$\frac{d\widetilde{u}_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}\widetilde{u}_{k}-\widetilde{a}_{1}\widetilde{u}_{j}+a_{2}\widetilde{u}_{j}\Phi_{j}\ast\bigg{(}\frac{u_{j}}{1+c_{u}u_{j}}\bigg{)},$ (7b)
$\frac{dv_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}v_{k}+v_{j}(b_{0}-b_{1}v_{j})-b_{3}\widetilde{u}_{j}v_{j}\widetilde{v}_{j}-\frac{b_{2}v_{j}}{1+c_{v}v_{j}}\Phi_{j}\ast\widetilde{v}_{j},$ (7c)
$\frac{d\widetilde{v}_{j}}{dt}$ $=-\sum_{k=1}^{N}L_{jk}\widetilde{v}_{k}-\widetilde{b}_{1}\widetilde{v}_{j}+b_{3}\widetilde{u}_{j}v_{j}\widetilde{v}_{j}+b_{2}\widetilde{v}_{j}\Phi_{j}\ast\bigg{(}\frac{v_{j}}{1+c_{v}v_{j}}\bigg{)},$ (7d)
where $u_{j}$ and $\widetilde{u}_{j}$ are the densities of healthy and toxic amyloid beta, and $v_{j}$ and $\widetilde{v}_{j}$ are the densities of healthy and toxic tau proteins at the node $j$ in a given graph. Here, $N$ is the total number of nodes and $L_{ij}$ are the entries of the Laplacian matrix for the given network. The convolution and the Laplacian of a graph are defined in thompson2020 ; pal2022sr .
There are nonlocal mathematical models on neural networks, especially convolutional neural networks. The traditional neural network blocks feature representations in a local sense; however, long-range dependencies have significant practical learning problems. Here, we describe a nonlocal network in the context of traditional image classification tasks and make a comparison among different neural networks. Suppose $X=[X_{1},\ldots,X_{M}]$ is an input sample with $X_{j}(j=1,\dots,M)$ as the feature at position $j$. A nonlocal output signal at position $j$ can be defined as tao2018 :
$Z_{j}=X_{j}+\frac{W_{Z}}{\mathcal{C}_{j}(X)}\sum_{\forall k}\omega(X_{j},X_{k})g(X_{k}),$
$Z^{n+1}:=Z^{n}+\mathcal{F}(Z^{n};W^{n}),\leavevmode\nobreak\ \leavevmode\nobreak\ n=0,1,\ldots,N,$
Genetic regulatory networks (GRNs) serve an important role in coordinating gene expression and orchestrating complex developmental processes in cells. These networks involve transcription factors binding to particular DNA sequences, cis-regulatory elements, and other substances, resulting in a dynamic and linked web of regulatory interactions. Understanding the abnormal alterations in GRNs gives insight into several disease pathways, such as cancer and developmental disorders nalls2021 . Stochastic transcriptional regulation adds a layer of complexity to GRNs, altering cellular heterogeneity, phenotypic variability, and the general adaptability of biological systems melnik2009a ; thieu2021 ; kim2023 . Furthermore, pattern memory in GRNs is important in cellular decision-making processes, allowing cells to respond to the same stimuli differently based on their past experience.
The spatio-temporal dynamics of biomolecular networks and biochemical systems relate to the complex, dynamic patterns of molecular interactions that occur both in space (inside cellular compartments) and over time kim2018 . The spatial organization impacts many biomolecular interactions that occur in various cellular compartments, such as the nucleus, cytoplasm, and organelles. Furthermore, spatiotemporal dynamics in metabolic networks are critical for coordinating energy generation, nutrient use, and maintaining cellular homeostasis peukert2014 . Nonetheless, spatial gradients of signalling molecules contribute to the generation of concentration gradients, impacting cell fate and pattern formation throughout development. The dynamics of transcriptional regulation involve both local and nonlocal components, reflecting the complex interplay of molecular processes within a cell as well as interactions that extend beyond immediate neighbours. Local dynamics often relate to nearby activities, such as interactions between genes and transcription factors. In contrast, nonlocal dynamics involve events that affect distant genomic areas or operate across distant locations and temporal dimensions.
@@ 2.6 Nonlocal dispersal
Nonlocal dispersal in biology refers to the movement or migration of individuals, cells, or organisms over extended distances, often beyond their immediate vicinity. This phenomenon plays a crucial role in shaping populations’ distribution, dynamics, and persistence in various ecological and biological systems. Species expand their range of interactions through dispersal, which has many forms. Till now, we have considered the simplest type of dispersal as random diffusion, which captures the movement of organisms between adjacent spatial locations. A different type of dispersal that gives rise to nonlocality is studied in the literature, where organisms can travel some distance li2014 . Periodic travelling wave solutions have been studied in prey-predator models with nonlocal dispersal to capture cyclic population dynamics. In sherratt2016 , the author considered the Rosenzweig-MacArthur model with nonlocal dispersal as:
@@
denotes the nonlocal dispersal corresponding to $u$ at a spatial point $x$ with $l$ as the spatial scale representing the dispersal distance li2014 ; $u(x,t)$ and $v(x,t)$ are the population densities of the prey and predator, respectively, at the spatial point $x$ and at time $t>0$. This type of nonlocal model is the derivative-free integral equation for diffusion khodabakhshi2021 . To study the oscillatory dynamics of the model, the author converted this model into normal form near the Hopf bifurcation using the complex Ginzburg-Landau equation garcia2012 . He estimated the wavelength and amplitude of the periodic travelling wave generated by the invasion of the coexisting steady-state for the parameter values close to the Hopf bifurcation threshold. Assuming specific properties for kernel function $\phi$ (non-negativity, being even, and normality), the nonlocal dispersal operator can be approximated for small $l$ as li2014 :
@@
On the other hand, for $l\rightarrow\infty$ and for a periodic $u$ with periodicity $L$, the operator $\mathcal{P}$ can be expressed as li2014 :
@@ 2.7 Fractional calculus and nonlocal models
Fractional nonlocal models have gained significant attention in biology due to their ability to capture the memory effect, anomalous transport, and long-range interactions, which are common in biological systems. Fractional calculus has been employed to study drug release from polymeric matrices used in drug delivery systems. These models account for drug diffusion’s nonlocal and memory effects within the polymer, providing a more accurate representation of drug release kinetics. Different nonlocal RD models have been used in biology and other fields. Researchers have used fractional nonlocal models to understand cancer cell populations’ dynamics in cancer diseases. Before going to the fractional model, we first consider a modified nonlocal RD model (3) for nonlocal consumption of resources as zhan2022 :
@@
The parameters $m$ and $n$ are positive integers, $m=1,n=1$ and $m=2,n=1$ correspond to the asexual and sexual reproductions, respectively. This model is an advanced version of the nonlocal model kinds of (3). Here, $a$, $b$ and $c$ are the parameters, and $\phi\ast u^{n}$ is the convolution. The last term in (9) represents the mortality with a rate constant $c$. In cancer disease, abnormal changes in cell density cause significant morbidity and mortality. Numerous mathematical models in the literature account for this disease propagation alarcon2003 . The fractional RD model is one of them, and many researchers have been focused on it. The time-fractional RD equation is used in the mathematical model to incorporate such an evolution process, and it is given by zhan2022 :
@@
Here, $u^{\prime}(\mathbf{x},s)$ is the first-order partial derivative of $u(\mathbf{x},t)$ with respect to $t$ and evaluated at $t=s$. These models have many applications, and some of them are mentioned in the next section.

### v2

@@ 2.2 Nonlocal interactions in a single species population
Without loss of our generality, we assume that $u$ is a dimensionless population density, and it is defined in the one-dimensional spatial domain $\Omega=[-L,L]\subset\mathbb{R}$, $L>0$. In the considered case $g(u)=u(1-u)$, the population has the reproduction term, which is proportional to its density and available resources $(1-u)$ [72]. For the local interaction, the population’s consumption rate at any given spatial point $x\in\Omega$ is proportional to its population’s density at the same spatial point with the total amount of available resources unity. In real-world systems, individuals do not always interact only with their immediate neighbours; instead, they may interact over longer distances due to long-range communication or interactions (e.g., chemical signalling in biological systems [82]), long-range dispersal (e.g., seeds blown by the wind or animals migrating over long distances [83]), long-range delay [84], and nonlocal competition or cooperation between species in ecology [47, 85, 86]. In all these cases, the individual located at a spatial point $x$ interacts with others in its neighbourhood locations [47, 86, 87, 88, 89, 90]. We assume that the individuals located at the spatial point $y$ access the resources of the point $x$ with the probability density function $\phi(x-y)$. In this case, the RD equation (2) modified into an integro-differential equation
@@
with non-negative initial and periodic boundary conditions over the domain $[-L,L]$. Here $\phi\ast u$ is the convolution of $u$, and we use `$\ast$` ∗’ as the convolution operator for the rest of this manuscript. The convolution $\phi\ast u$ in one spatial dimension is defined by
@@
This kernel function represents how individuals at position $y$ influence those at position $x$. For instance, if the kernel is Gaussian, it represents the strongest interactions at close distances but tapers off smoothly. In physics, it defines systems in which individuals can influence others locally and globally through long-range interactions. In addition, the RD equation (2) fails to incorporate such long-range interactions [90]. The nonlocal model (3) has two homogeneous stationary solutions $u=0$ and $u=1$. In the case of the local model, the solution $u=1$ is stable under heterogeneous perturbations, but it may not be stable for the nonlocal model [72]. A brief analysis of this single species nonlocal model is given in the next section.
@@ 2.3 Nonlocal interactions in time and space
Researchers have studied nonlocal RD models for the triangular, parabolic kernel functions to account for different biological phenomena [91, 92, 93, 94, 54, 95, 96, 97, 98]. These kernel functions are known as the truncated kernel. Such non-truncated kernel functions have been the subject of intensive studies. For instance, Britton introduced an advanced version of the single species model (3) as:
@@
where $\alpha u$ ($\alpha>0$) represents the local aggregation and $-(1+\alpha)\psi\ast u$ is the disadvantage in the global population as the resources may go into depletion of being high population density [99]. The author studied the nonlocal model with the Laplacian kernel function [54]. Sometimes, species take time to move; in this case, the correct average population density is a spatio-temporal weighted average toward the current time and position. Taking this spatio-temporal factor into account, Britton introduced a space-time nonlocality in the single species model (4) as [100]:
@@
Likewise, the spatial kernel function $\phi$ in (3), the spatio-temporal kernel function $\psi$ satisfies some conditions mentioned in [100, 101]. In some cases, these types of nonlocal models with non-truncated kernels can be handled by converting them into coupled RD equations [99, 100]. However, in many other cases, this may not be possible. Indeed, nonlocal time-dependent interaction occurs frequently in physics and engineering, with nonlocal time affecting the dynamics of open quantum systems [102, 103], stress depending on the current strain and its history [104], along with many other situations where nonlocality has to be accounted for.
@@ 2.4 Nonlocal interactions in cell biology and genomics
In cell biology, nonlocal interactions are interactions between cellular components that extend beyond immediate neighbours and involve long-distance signalling or communication. For instance, mechanical forces inside tissues influence cell migration and division over considerable distances from the point of force application, as tension and compression are conveyed through the cytoskeleton and extracellular matrix [105]. These forces travel like waves or through elastic deformation. The transmission of forces follows principles of elasticity, where cells act like soft materials that deform under force. Additionally, tumors are abnormal tissue masses of cells with different phenotypes, such as cancer stem cells [106, 107, 108], and they are heterogeneously distributed in the body. Therefore, nonlocal interactions are crucial in cellular processes and behaviours such as cell migration, pattern formation, and collective cell responses. Motivated by the paradoxical findings of an agent-based model [109], Hillen et al. have proposed a mathematical model [110]:
@@
Gene silencing and RNA interference are exciting fields of study that have promising implications in health and industry [111]. For instance, a synthetic RNA molecule is designed to selectively target and mute the production of one particular gene in small interfering RNA technology. When it enters cells, it attaches itself to the target mRNA, causing it to be degraded and stopping the creation of new proteins. There are some more key technologies and models associated with gene silencing and RNA interference: short hairpin RNA [112], micro RNA [113], CRISPR/Cas9 technology [114], Antisense Oligonucleotides [115], Dicer Knockout models [116], etc. It is essential to model genetic switches and networks to comprehend the complex regulatory processes within cells [117]. The dynamics of genetic switches and networks have been simulated and examined using various computational and mathematical models. Some common approaches are Boolean networks, ordinary differential equations, stochastic models, agent-based models, and machine learning approaches. Furthermore, nonlocal models are important beyond DNA, RNA, and protein modelling. The use of microarray technology for placing the full human genome on chips is one such application.
@@
A multi-scale modelling technique is required to represent a cell as a composite structure incorporating microtubules and organelles. Several models have been created to simulate these components’ interactions and dynamics, such as continuum mechanics models [118, 119, 120]. Understanding the methods by which a cell develops, duplicates its components, and divides into two daughter cells is fundamental to studying cell cycles. Modelling cell cycles with fractional derivatives presents a mathematical technique reflecting cellular dynamics’ nonlocal and memory-dependent character. In addition, rheology and hysteresis are more frequently linked with materials and physical systems, and they may also be used to describe some aspects of cellular activity, including cell mechanics and cycles [121]. Understanding the interaction of grain shape, material characteristics, and nonlocal rheology is critical for forecasting and regulating the behaviour of dense granular flows [122].
@@
Different types of biological rhythms occur in various physiological and behavioural processes in living organisms, and they are essential for maintaining homeostasis and optimizing physiological functions in organisms. In addition, angiogenesis is the physiological process that results in forming new blood vessels from pre-existing vessels [123]. This complex mechanism is required for a variety of physiological activities, including embryonic development and wound healing. However, angiogenesis has been linked to several pathological disorders, including cancer and inflammatory illnesses. Furthermore, biological cells are complex structures with distinct rheological properties, and understanding their mechanical responses is critical for a variety of applications, including disease diagnostics, medication delivery, and tissue creation. Studying these rhythms provides insights into how living organisms adapt to their environments and regulate various processes over time.
@@ 2.5 Nonlocal dispersal
Nonlocal dispersal in biology refers to the movement or migration of individuals, cells, or organisms over extended distances, often beyond their immediate vicinity. This phenomenon plays a crucial role in shaping populations’ distribution, dynamics, and persistence in various ecological and biological systems. Species expand their range of interactions through dispersal, which has many forms. Till now, we have considered the simplest type of dispersal as random diffusion, which captures the movement of organisms between adjacent spatial locations. A different type of dispersal that gives rise to nonlocality is studied in the literature, where organisms can travel some distance [124]. In [54], the author considered the Rosenzweig-MacArthur model with nonlocal dispersal as:
@@
denotes the nonlocal dispersal corresponding to $u$ at a spatial point $x$ with $l$ as the spatial scale representing the dispersal distance [124]; $u(x,t)$ and $v(x,t)$ are the population densities of the prey and predator, respectively, at the spatial point $x$ and at time $t>0$. This type of nonlocal model is the derivative-free integral equation for diffusion [125]. To study the oscillatory dynamics of the model, the author converted this model into normal form near the Hopf bifurcation using the complex Ginzburg-Landau equation [126]. He estimated the wavelength and amplitude of the periodic travelling wave generated by the invasion of the coexisting steady-state for the parameter values close to the Hopf bifurcation threshold. Assuming specific properties for kernel function $\phi$ (non-negativity, being even, and normality), the nonlocal dispersal operator can be approximated for small $l$ as [124]:
@@
On the other hand, for $l\rightarrow\infty$ and for a periodic $u$ with periodicity $L$, the operator $\mathcal{P}$ can be expressed as [124]:
@@
In conventional RD models, particle flux is typically determined by local concentration gradients (Fick’s laws). However, the nonlocal dispersal is a generalization of random walks where the particles can jump to further locations instead of stepping to nearby sites [127]. This type of dispersal can be facilitated by various natural phenomena, including wind and water currents. For instance, seeds may be blown across great distances by the wind or migratory birds.
@@ 2.6 Fractional calculus and nonlocal models
Fractional nonlocal models have gained significant attention in biology due to their ability to capture the memory effect, anomalous transport, and long-range interactions, which are common in biological systems. Fractional calculus has been employed to study drug release from polymeric matrices used in drug delivery systems. These models account for drug diffusion’s nonlocal and memory effects within the polymer, providing a more accurate representation of drug release kinetics. Different nonlocal RD models have been used in biology and other fields. Researchers have used fractional nonlocal models to understand cancer cell populations’ dynamics in cancer diseases. Before going to the fractional model, we first consider a modified nonlocal RD model (3) for nonlocal consumption of resources as [128, 97, 98]:
@@
The parameters $m$ and $n$ are positive integers, $m=1,n=1$ and $m=2,n=1$ correspond to the asexual and sexual reproductions, respectively. This model is an advanced version of the nonlocal model kinds of (3). Here, $a$, $b$ and $c$ are the parameters, and $\phi\ast u^{n}$ is the convolution. The last term in (8) represents the mortality with a rate constant $c$. In cancer disease, abnormal changes in cell density cause significant morbidity and mortality. Numerous mathematical models in the literature account for this disease propagation [129]. The fractional RD model is one of them, and many researchers have been focused on it. The time-fractional RD equation is used in the mathematical model to incorporate such an evolution process, and it is given by [128]:
@@
Here, $u^{\prime}(\mathbf{x},s)$ is the first-order partial derivative of $u(\mathbf{x},t)$ with respect to $t$ and evaluated at $t=s$. This type of time-fractional model incorporates memory effects where the current state of a system depends not only on its local conditions but also on its historical states [130, 131]. In addition, fractional nonlocal models are used to describe anomalous diffusion, where the spread of particles does not follow classical Brownian motion [132]. The key feature of anomalous diffusion is that the mean square displacement scales as a nonlinear power law in time ($\delta x^{2}\right>=t^{\beta}\right)$. Superdiffusion refers to faster than linear scaling, i.e., $\beta>1$, while subdiffusion refers to slower than linear scaling, i.e., $\beta<1$ [133].